Circular No. 8847

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 133P/ELST-PIZARRO D. Jewitt, P. Lacerda, and N. Peixinho, University of Hawaii, report

that comet 133P/Elst-Pizarro = minor planet (7968) Elst-Pizarro has become active after a long period of quiescence. Optical observations with the University of Hawaii 2.2-m telescope on June 11 UT show a straight tail at least 20′′ long in p.a. 256^{o}. The apparent red magnitude within an aperture 8′′ in projected diameter is ≈ 19.5. Activity in 133P was last observed in 2002 December. The re-emergence of activity near perihelion (T = 2007 June 29.3 TT; e.g., Nakano 2007, ICQ 2007 Comet Handbook) is consistent with the identification of this object as an ice-bearing minor planet or “main-belt comet” (Hsieh et al. 2004, A.J. 127, 2997). Further observations to characterize the evolution of the mass loss in the coming months are encouraged.

COMETS C/2007 F2–F5, C/2007 G2, C/2007 H4–H9 (SOHO) Additional near-sun comets (cf. IAUC 8846) have been found on SOHO

website images — all being Kreutz sungrazers except for C/2007 F4 (Meyer group). C/2007 F2 reached mag ∼ 5 in C3 images, and showed a thin, faint, 15′ tail in C2 images. C/2007 F3 was stellar in appearance, reaching mag ∼ 6; it was tailless even in C2 images. C/2007 F4 was stellar in appearance and reached mag ∼ 6.5. C/2007 H6, C/2007 H8, and C/2007 H9 were fuzzy, very faint, and tailless. C/2007 H7 was was slightly diffuse, reaching mag ∼ 7, with a hint of a tail. The remaining four objects were very small, somewhat diffuse, and quite faint (mag ∼ 7.5–8).

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 F2 Mar. 20.904 0^{h}29.^{m}3 − 1^{o}36′ C3/2 BZ 2007-K68
C/2007 F3  27.446 0 40.2 + 0 46 C3/2 BZ 2007-K68
C/2007 F4  28.488 0 32.0 + 4 45 C3/2 BZ 2007-K69
C/2007 F5 30.118 0 42.0 + 3 08 C2 TH 2007-K69
C/2007 G2 Apr. 8.018 1 14.4 + 6 19 C2  TH 2007-K69
C/2007 H4  18.963 1 53.4 +10 03 C2  BZ 2007-K69
C/2007 H5  20.743 1 59.9 +10 36 C2  RK 2007-K69
C/2007 H6  21.868 2 04.5 +11 04 C2  VB 2007-L02
C/2007 H7  23.160 2 08.7 +11 18 C2  BZ 2007-L02
C/2007 H8  29.976 2 33.0 +13 26 C2  BZ 2007-L02
C/2007 H9  30.535 2 35.6 +13 32 C2  BZ 2007-L02
2007 June 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green